. What is the ratio of nuclear densities of two nuclei having mass numbers in the ratio 1:8? 


1
Expert's answer
2021-04-13T06:30:38-0400

We have:


"\frac{m_1}{m_2}=\frac{\rho_1 V_1}{\rho_2V_2}=\frac18."

re-write in terms of radii:

"\frac{\rho_1r_1^3}{\rho_2r_2^3}=\frac{\rho_1[R_0(m_1)]^{1/3}}{\rho_2[R_0(m_2)]^{1/3}}=\frac18,\\\space\\
\frac{\rho_1·1^{1/3}}{\rho_2·8^{1/3}}=\frac18,\\\space\\
\frac{\rho_1}{\rho_2}=\frac14."